Fay’s uses a perpetual inventory system. In March, Fay’s purchased a type of swimwear designed to be slimming to the wearer. It purchased twenty suits of varying sizes for $40 each and priced them at $120 each. They sold out almost immediately, so Fay purchased forty more suits in April for $40 each and sold thirty-eight of them for $130 each. Again in July, Fay made one more purchase of twenty suits at $40 each and sold fifteen of them for $130 each. Fay decided not to put the rest of her inventory on sale at the end of the summer, but to hold onto it until cruise season started the following winter. She believed she could sell the rest then without having to mark them down.

1.

Make the journal entries for the purchases Fay made.

2.

Make the journal entries for the sales Fay made.

3.

Determine the balance in ending inventory on December 31.



1
Expert's answer
2021-07-06T12:09:39-0400

Solution:

1.). Journal entries for the purchases Fay made are as follows:



 

 

2.). Journal entries for the sales Fay made are as follows:



 

3.). The balance of the ending inventory = 27 suits.

Total inventory purchased = 20 + 40 + 20 = 100

Total inventory sold = 20 + 38 + 15          =  73


Total inventory balance  (100 - 73)               = 27